OpenStudy (anonymous):

Find the equation of the tangent to the curve defined by y = x + ln x that is perpendicular to the line defined by 3x + 9y = 8

OpenStudy (anonymous):

First we want to get the slope which we will get from this equation 3x + 9y = 8 Can you tell me what the slope of this line is?

OpenStudy (anonymous):

i just need the steps and the answer. i already learned the material back in september, just need some refreshing for the exam i have next week.

OpenStudy (anonymous):

Oh ok First get the slope. Then find the derivative of x + ln x Set the derivative equal to the slope and solve for x. Plug the x into the original equation y = x + ln x to get y Now you have your slope m and point (x,y) Use point slope form (y-y1)=m(x-x1) where x1 and y1 is your point

OpenStudy (anonymous):

You can leave it as it is or simplify it to y-intercept form.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

word, now i remember. thanks.

OpenStudy (anonymous):

The slope is the negative reciprocal of the slope of this line 3x + 9y = 8 just fyi